OpenStudy (anonymous):

What is the sum of the multiples of 3 between 3 and 999, inclusive?

OpenStudy (anonymous):

n = 333 \[\large Sum = \frac{333}{2}(3 + 999)\]

OpenStudy (anonymous):

3n = 999 n = 333...

OpenStudy (anonymous):

for the first thing you told me to solev i got 5511

OpenStudy (anonymous):

\[\large Sum = \frac{333}{2}(3 + 999)\]

OpenStudy (anonymous):

and i got 166,833 for the second one

OpenStudy (anonymous):

Solve this thing..

OpenStudy (anonymous):

Yes second one is right..

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

3 + 6 + 9 + ........................+ 999 a = 3 d= 3 and last term is 999: 999 = 3 + (n-1)(3) 999 = 3n n = 333.. So use this: \(\large Sum=\frac{333}{2}(3+999)\)

OpenStudy (anonymous):

Why?? Do you have any doubt in that??

OpenStudy (anonymous):

no.... i have to know the answer obviously thats why

OpenStudy (anonymous):

before i put it down and its wrong

OpenStudy (anonymous):

That is the right answer..

OpenStudy (anonymous):

ohk great! thanks so much!

OpenStudy (anonymous):

Welcome..